# Boundary conditions matter: On the spectrum of infinite quantum graphs

### v2

@@ Abstract.
The work of J.B.K. was supported by the Fundação para a Ciência e a Tecnologia, Portugal, via project PTDC/MAT-PUR/1788/2020. The work of D.M. and M.P. was supported by the Deutsche Forschungsgemeinschaft (Grant 397230547). This article is based upon work from COST Action 18232 MAT-DYN-NET, supported by COST (European Cooperation in Science and Technology), www.cost.eu.
@@ 1. Introduction
Infinite quantum graphs, and in particular their spectral properties, have been studied since the late 1990s in [13, 15, 7, 12] and more recently among others in [42, 5]. Therein, the relevant operator is essentially self-adjoint. The case where different self-adjoint extensions may exist was first discussed in the pioneering papers [14, 50], and then more extensively in [21, 38]. In these papers, the spectral properties of only the Friedrichs extension of the (general realisations of) Laplacians on infinite quantum graphs were discussed explicitly. First investigations relied on a transference principle relating the spectrum of a quantum graph to the spectrum of an infinite matrix: this idea goes back to von Below’s early analysis in [6] and relies fundamentally upon the assumption that all edges have the same length. If this assumption is dropped, infinite quantum graphs will generally not be essentially self-adjoint: the operator domain may then have to include appropriate boundary conditions at infinity, which in turn will influence the spectrum. A similar phenomenon is known for discrete Laplacians on combinatorial graphs, see e.g. [31, 23].
@@
As intimated above, the notion of infinite quantum graph we are going to discuss relies on the notion of ends, a classical concept from graph theory (see, e.g., [19, Chapter 8] and references therein) which was recently adapted to quantum graphs in [37]. As shown therein, if an end is “thin enough” in a certain sense, then different boundary conditions can be imposed on it leading to different self-adjoint realisations of the Laplacian. We will treat what are arguably the most important two types of conditions, Dirichlet and Neumann, in this paper.
@@
Let us also point out parallels of this paper to the spectral geometry on domains and manifolds – a well developed field, see, e.g. [25] for a survey. If $\Omega\subset\mathbb{R}^{d}$ is a bounded domain with Lipschitz boundary, then its Dirichlet and Neumann Laplacians both have purely discrete spectrum, whereas on unbounded or non-smooth domains, the situation is more subtle, at least in the Neumann case, see, e.g., [1, Chapter 6] or [18]. In particular, a family of bounded planar domains (the combs) was introduced in [29] that displays a remarkable behaviour: the Neumann Laplacian thereon has empty absolutely continuous spectrum but its essential spectrum is nonempty and contains infinitely many dilated copies of the set of accumulation points of some (possibly unbounded) sequence, see [29, Theorem 3.4].
@@
In the case of infinite metric graphs, the issue of discreteness of the spectrum is significantly less well understood: examples are known that show that the spectrum may or may not be purely discrete [59, 26, 38, 24]. Remarkably, even infinite equilateral graphs may have purely discrete spectrum: this can, e.g., be seen by applying a well-known transference principle [15] to suitable infinite combinatorial graphs with purely discrete spectrum, like the infamous lamplighter graph (all of whose eigenvalues have infinite multiplicity). There already exist some abstract criteria for the compactness of embeddings of Sobolev spaces over metric measure spaces, as introduced in [27, 17] see, e.g., [17, Theorem 4.59] and [33, Theorem 2]. In particular, it has been known since [14] that on metric graphs the spectrum is discrete provided the graph has finite volume. Other criteria of a more geometric flavour have been provided in [38, Corollary 3.5 and Corollary 4.5] and [21, Theorem 4.18]. We also mention [59], which provides criteria for the discreteness of the spectrum and investigates the asymptotics of the eigenvalues on so-called regular trees.
@@
This paper is organised as follows: Section 2 contains definitions of metric graphs, Sobolev spaces, and extensions of the Laplace operator. In Section 3, we provide criteria on discreteness or non-discreteness of the spectrum of these Laplacian extensions: our main contribution in this context is Theorem 3.1, a new Kolmogorov–Riesz-type criterion that, in particular, gives a sufficient and necessary condition for compactness of the embedding of Sobolev spaces over metric graphs $\mathcal{G}$ into the canonical $L^{2}$-space in terms of “thinness” of the periphery of the graph, measured by the $L^{2}$-norm of functions supported outside compact subgraphs of $\mathcal{G}$. Another criterion for discreteness of the Friedrichs extension on infinite trees from [38] is shown not to generalise to the Neumann extension. Subsection 3.2 puts these criteria to use by investigating a parameter-dependent family $\mathcal{G}_{\alpha}$ of metric graphs (the diagonal combs) on which the spectra of the Neumann extensions exhibit a noteworthy phase transition from purely discrete to nonempty essential spectrum, despite all having infinite volume: a phenomenon which somewhat recalls Neumann Laplacians on (finite volume) cusp domains but which has no parallels on domains of infinite volume. To the best of our knowledge, Theorem 3.1 is the first criterion able to identify this transition in the context of quantum graphs. Beyond quantum graphs, similar phenomena are elusive and have been discovered only recently: we mention e.g. the celebrated results about phase transition from pure point spectrum to absolute continuous spectrum for almost Mathieu operators, see [32] and references therein.
@@ 2.1. Notation and elementary properties
A metric graph is indeed a metric space with respect to the canonical metric induced by the one-dimensional Euclidean distance and by the combinatorial distance on the underlying discrete graph, see [49] for details.
@@
$\operatorname{Diam}(\mathcal{G}):=\sup_{x,y\in\mathcal{G}}\operatorname{dist}_{\mathcal{G}}(x,y),$
@@
A metric graph is also, in a canonical way, a measure space, endowed with the direct sum of Lebesgue measures on each interval $(0,\ell_{\mathsf{e}})$. Throughout the article we will assume that $\mathcal{G}$ is locally finite and connected, which implies in particular that $\mathcal{G}$ is a metric measure space, cf. [60, Section 3], in the sense that each $x\in\mathcal{G}$ has a neighbourhood of finite measure.
@@
    We call $\mathcal{G}$ a tree if it contains no cycles as induced subgraphs.
@@
Whenever $\mathsf{E}$ is infinite, the ends of $\mathcal{G}$ can be defined as in [37, Definition 3.1] based on a classical combinatorial approach. For our purposes, a different (but equivalent, see [37, Section 2.2] and references therein) notion of end of topological flavour is more convenient.
@@
Consider a sequence $\mathcal{U}=(U_{n})$ of nonempty open connected subsets of $\mathcal{G}$ with compact boundaries and such that $U_{n+1}\subset U_{n}$ for all $n\in\mathbb{N}$ and $\bigcap_{n\in\mathbb{N}}\overline{U_{n}}=\emptyset$. Two such sequences $\mathcal{U}=(U_{n})$ and $\mathcal{U^{\prime}}=(U^{\prime}_{n})$ are called equivalent if for all $n\in\mathbb{N}$ there exist $j,k$, such that $U^{\prime}_{j}\subset U_{n}$ and $U_{k}\subset U^{\prime}_{n}$. An equivalence class $\gamma$ of sequences is called a topological end – or simply an end – of $\mathcal{G}$ and $\mathfrak{C}(\mathcal{G})$ denotes the set of all (topological) ends of $\mathcal{G}$.
@@
At the risk of redundancy, let us stress that ends are not vertices. In particular, an end cannot be an endpoint of an edge. For this reason a subset or subgraph of $\mathcal{G}$ is compact if and only if it is closed in $\mathcal{G}$ and intersects a finite number of edges of $\mathcal{G}$.
@@
[…]
@@ 2.2. Compact exhaustions
For an infinite graph $\mathcal{G}$, we are going to consider appropriate sequences of compact subgraphs $\mathcal{G}_{n}$ which approximate $\mathcal{G}$ in a suitable way (cf. [37, Section 2.2]). This will allow us to generalise many properties of $\mathcal{G}_{n}$ directly to $\mathcal{G}$, in particular some eigenvalue bounds.
@@
We will often use the following construction, which yields a compact exhaustion in any locally finite, connected graph $\mathcal{G}$: Fix a vertex $\mathsf{v}\in\mathsf{V}$. Then, any $\mathsf{w}\in\mathsf{V}$ has a well-defined combinatorial distance to $\mathsf{v}$, see for instance [19, Section 1.3]. Set
@@
$\operatorname{Diam}(\mathcal{G})\leq\liminf_{n\to\infty}\operatorname{Diam}(\mathcal{G}_{n}).$
@@ 2.3. Lebesgue and Sobolev spaces
We next introduce function spaces on a metric graph $\mathcal{G}$, for which embedding theorems will be discussed in Section 3. Any metric graph $\mathcal{G}$ has both a topological and a measure theoretical structure. This immediately defines the space of continuous functions and Lebesgue spaces; we refer to [49] and [37, Sections 2.3 and 3.1] for the details of these function spaces.
@@
$L_{c}^{2}(\mathcal{G}):=\left\{f\in L^{2}(\mathcal{G}):f(x)=0\text{ almost everywhere outside a compact subset of }\mathcal{G}\right\}.$
@@
Replacing $L^{2}$-spaces by $L^{p}$-spaces in the definition, we also obtain the Sobolev spaces $W^{1,p}(\mathcal{G})$ for $p\geq 1$; in particular, $W^{1,2}(\mathcal{G})=H^{1}(\mathcal{G})$. It is not hard to show that, for any $f\in H^{1}(\mathcal{G})$, there is a canonical extension of $f$ to the ends of $\mathcal{G}$ (see [37, Definition 3.4]). In particular, if $\overline{\mathcal{G}}$ is compact (see Corollary A.11 for a sufficient condition), then by [37, Lemma 3.2] $f\in H^{1}(\mathcal{G})$ can be identified with a function on $\overline{\mathcal{G}}$, and up to a canonical identification, we have $H^{1}(\mathcal{G})\subset C(\overline{\mathcal{G}})$.
@@
$H^{1}_{c}(\mathcal{G}):=\left\{f\in H^{1}(\mathcal{G}):f(x)=0\text{ outside a compact subset of }\mathcal{G}\right\}$
@@
$H^{1}_{0}(\mathcal{G}):=\overline{H^{1}_{c}(\mathcal{G})}^{\|\cdot\|_{H^{1}(\mathcal{G})}}.$
@@
    that is, $H^{1}_{0}(\mathcal{G})$-functions satisfy a Dirichlet (zero) condition at every end, see [37, Theorem 3.12].
@@
    A unified way to prescribe Dirichlet conditions at vertices, is as follows: remove $\mathsf{v}$ and appropriately add infinitely many dummy vertices on all edges incident to it. This way, $\mathsf{v}$ is replaced by $\deg(\mathsf{v})$ many ends $\gamma$, and every $f\in H^{1}_{0}(\mathcal{G})$ must vanish at $\mathsf{v}$. While being technical, this approach allows to refer to and apply a result from [38] in our setting, see Proposition 3.2.

### v3

# On the spectrum of infinite quantum graphs

@@ Abstract.
The work of J.B.K. was supported by the Fundação para a Ciência e a Tecnologia, Portugal, via projects 2023.13921.PEX, https://doi.org/10.54499/2023.13921.PEX, UIDB/00208/2020, https://doi.org/10.54499/UIDB/00208/2020 and UIDB/04106/2020, https://doi.org/10.54499/UIDB/04106/2020. The work of D.M. and M.P. was supported by the Deutsche Forschungsgemeinschaft (Grant 397230547). This article is based upon work from COST Action 18232 MAT-DYN-NET, supported by COST (European Cooperation in Science and Technology), www.cost.eu.
@@ 1. Introduction
Infinite quantum graphs, and in particular their spectral properties, have been studied since the late 1990s in [18, 20, 12, 17] and more recently among others in [47, 10]. Therein, the relevant operator is essentially self-adjoint. The case where different self-adjoint extensions may exist was first discussed in the pioneering papers [19, 54], and then more extensively in [13, 25, 44]. In these papers, the spectral properties of only the Friedrichs extension of the (general realisations of) Laplacians on infinite quantum graphs were discussed explicitly. First investigations relied on a transference principle relating the spectrum of a quantum graph to the spectrum of an infinite matrix: this idea goes back to von Below’s early analysis in [11] and relies fundamentally upon the assumption that all edges have the same length. If this assumption is dropped, infinite quantum graphs will generally not be essentially self-adjoint: the operator domain may then have to include appropriate boundary conditions at infinity, which in turn will influence the spectrum. A similar phenomenon is known for discrete Laplacians on combinatorial graphs, see e.g. [36, 28]. Also, let us point out that it has been known since [7], and implicitly since [5], that certain classes of fractals, including so-called Hanoi attractors, can be approximated by metric graphs: in other words, Laplacians thereon are infinite quantum graphs, in our language. Spectral analysis for such classes of fractals has been performed since [4, 6, 3]. However, we do note that important classes of infinite quantum graphs, even as natural as the Neumann realisation of the Laplacians on (fractal-like, but possibly non-selfsimilar) radially symmetric trees with infinitely many ends, are not covered by the theory in the present paper; and yet, it is known that under suitable growth conditions, they may have pure point spectrum: we refer to [62, 56, 38] and especially to the recent paper [26].
@@
As intimated above, the notion of infinite quantum graph we are going to discuss relies on the notion of ends, a classical concept from graph theory (see, e.g., [23, Chapter 8] and references therein) which was recently adapted to quantum graphs in [43]. As shown therein, if an end is “thin enough” in a certain sense, then different boundary conditions can be imposed on it leading to different self-adjoint realisations of the Laplacian. We will treat what are arguably the most important two types of conditions, Dirichlet and Neumann, in this paper.
@@
Let us also point out parallels of this paper to the spectral geometry on domains and manifolds – a well developed field, see, e.g. [30] for a survey. If $\Omega\subset\mathbb{R}^{d}$ is a bounded domain with Lipschitz boundary, then its Dirichlet and Neumann Laplacians both have purely discrete spectrum, whereas on unbounded or non-smooth domains, the situation is more subtle, at least in the Neumann case, see, e.g., [1, Chapter 6] or [22]. In particular, a family of bounded planar domains (the combs) was introduced in [34] that displays a remarkable behaviour: the Neumann Laplacian thereon has empty absolutely continuous spectrum but its essential spectrum is non-empty and contains infinitely many dilated copies of the set of accumulation points of some (possibly unbounded) sequence, see [34, Theorem 3.4].
@@
In the case of infinite metric graphs, the issue of discreteness of the spectrum is significantly less well understood: examples are known that show that the spectrum may or may not be purely discrete [62, 31, 44, 29]. Remarkably, even infinite equilateral graphs may have purely discrete spectrum: this can, e.g., be seen by applying a well-known transference principle [20] to suitable infinite combinatorial graphs with purely discrete spectrum, like the infamous lamplighter graph (all of whose eigenvalues have infinite multiplicity). There already exist some abstract criteria for the compactness of embeddings of Sobolev spaces over metric measure spaces, as introduced in [32, 21] see, e.g., [21, Theorem 4.59] and [39, Theorem 2]. In particular, it has been known since [19] that on metric graphs the spectrum is discrete provided the graph has finite volume. Other criteria of a more geometric flavour have been provided in [44, Corollary 3.5 and Corollary 4.5] and [25, Theorem 4.18]. We also mention [62], which provides criteria for the discreteness of the spectrum and investigates the asymptotics of the eigenvalues on so-called regular trees.
@@
This paper is organised as follows: Section 2 contains definitions of metric graphs, Sobolev spaces, and extensions of the Laplace operator. In Section 3, we provide criteria on discreteness or non-discreteness of the spectrum of these Laplacian extensions: our main contribution in this context is Theorem 3.1, a new Kolmogorov–Riesz-type criterion that, in particular, gives a sufficient and necessary condition for compactness of the embedding of Sobolev spaces over metric graphs $\mathcal{G}$ into the canonical $L^{2}$-space in terms of “thinness” of the periphery of the graph, measured by the $L^{2}$-norm of functions supported outside compact subgraphs of $\mathcal{G}$. Another criterion for discreteness of the Friedrichs extension on infinite trees from [44] is shown not to generalise to the Neumann extension. Subsection 3.2 puts these criteria to use by investigating a parameter-dependent family $\mathcal{G}_{\alpha}$ of metric graphs (the diagonal combs) on which the spectra of the Neumann extensions exhibit a noteworthy phase transition from purely discrete to non-empty essential spectrum, despite all having infinite volume: a phenomenon which somewhat recalls Neumann Laplacians on (finite volume) cusp domains but which has no parallels on domains of infinite volume. To the best of our knowledge, Theorem 3.1 is the first criterion able to identify this transition in the context of quantum graphs. Beyond quantum graphs, similar phenomena are elusive and have been discovered only recently: we mention e.g. the celebrated results about phase transition from pure point spectrum to absolute continuous spectrum for almost Mathieu operators, see [37] and references therein. Weaker but comparable results were recently obtained in the context of metric graphs in [45, 42].
@@ 2.1. Notation and elementary properties
A metric graph is indeed a metric space with respect to the canonical shortest-path metric induced by the one-dimensional Euclidean distance and by the combinatorial distance on the underlying discrete graph, see [53] for details.
@@
(2.1) $\operatorname{Diam}(\mathcal{G}):=\sup_{x,y\in\mathcal{G}}\operatorname{dist}_{\mathcal{G}}(x,y),$
@@
A metric graph is also, in a canonical way, a measure space, endowed with the direct sum of Lebesgue measures on each interval $(0,\ell_{\mathsf{e}})$. Throughout the article we will assume that $\mathcal{G}$ is locally finite and connected, which implies in particular that $\mathcal{G}$ is a metric measure space, cf. [63, Section 3], in the sense that each $x\in\mathcal{G}$ has a neighbourhood of finite measure.
@@
    We call $\mathcal{G}$ a tree if it contains no cycles as subgraphs.
@@
Whenever $\mathsf{E}$ is infinite, the ends of $\mathcal{G}$ can be defined as in [43, Definition 3.1] based on a classical combinatorial approach. For our purposes, a different (but equivalent, see [43, Section 2.2] and references therein) notion of end of topological flavour is more convenient.
@@
Consider a sequence $\mathcal{U}=(U_{n})$ of non-empty open connected subsets of $\mathcal{G}$ with compact boundaries and such that $U_{n+1}\subset U_{n}$ for all $n\in\mathbb{N}$ and $\bigcap_{n\in\mathbb{N}}\overline{U_{n}}=\emptyset$. Two such sequences $\mathcal{U}=(U_{n})$ and $\mathcal{U^{\prime}}=(U^{\prime}_{n})$ are called equivalent if for all $n\in\mathbb{N}$ there exist $j,k$, such that $U^{\prime}_{j}\subset U_{n}$ and $U_{k}\subset U^{\prime}_{n}$. An equivalence class $\gamma$ of sequences is called a topological end – or simply an end – of $\mathcal{G}$ and $\mathfrak{C}(\mathcal{G})$ denotes the set of all (topological) ends of $\mathcal{G}$.
@@
Let us stress that ends are not vertices. In particular, an end cannot be an endpoint of an edge. Thus a subset or subgraph of $\mathcal{G}$ is compact if and only if it is closed in $\mathcal{G}$ and intersects a finite number of edges of $\mathcal{G}$.
@@
A related concept to the metric completion $\overline{\mathcal{G}}$ is the Freudenthal compactification of $\mathcal{G}$, the coarsest extension of the canonical topology of $\mathcal{G}$ to $\mathcal{G}\cup\mathfrak{C}(\mathcal{G})$ which renders this space compact. The relation between these two spaces is a delicate business. We refer to Appendix A for a proper discussion and only summarise the following important facts: There is a canonical way to identify points in $\overline{\mathcal{G}}\backslash\mathcal{G}$ with ends $\gamma\in\mathfrak{C}(\mathcal{G})$, see Proposition A.1. However this map neither needs to be injective nor surjective, see Example A.7. Nevertheless, in important special cases, this map will lead to a homeomorphism which ensures that under certain conditions the notions of metric completion and the Freudenthal compactification coincide. This includes graphs where all ends have finite volume, cf. Corollary A.11 (hence in particular graphs of finite volume as in Sections 4 and 5), as well as graphs of finite diameter with only one end such as the diagonal combs investigated in Section 3.2.
@@ 2.2. Compact exhaustions
For an infinite graph $\mathcal{G}$, we are going to consider appropriate sequences of compact subgraphs $\mathcal{G}_{n}$ which approximate $\mathcal{G}$ in a suitable way (cf. [43, Section 2.2]). This will allow us to generalise many properties of $\mathcal{G}_{n}$ directly to $\mathcal{G}$, in particular some eigenvalue bounds.
@@
We will often use the following construction, which yields a compact exhaustion in any locally finite, connected graph $\mathcal{G}$: Fix a vertex $\mathsf{v}\in\mathsf{V}$. Then, any $\mathsf{w}\in\mathsf{V}$ has a well-defined combinatorial distance to $\mathsf{v}$, see for instance [23, Section 1.3]. Set
@@
(2.4) $\operatorname{Diam}(\mathcal{G})\leq\liminf_{n\to\infty}\operatorname{Diam}(\mathcal{G}_{n}).$
@@
Example 2.12.
We consider the infinite ladder graph $\mathcal{G}$ as depicted in Figure 2.1. The graph edges will be denoted by $\mathsf{e}_{n},\mathsf{f}_{n}^{1}$ and $\mathsf{f}_{n}^{2}$ for $n\in\mathbb{N}$ (see Figure 2.1 for the notation).
$\mathsf{f}_{1}^{1}$$\mathsf{f}_{2}^{1}$$\mathsf{f}_{3}^{1}$$\mathsf{f}_{4}^{1}$$\mathsf{f}_{1}^{2}$$\mathsf{f}_{2}^{2}$$\mathsf{f}_{3}^{2}$$\mathsf{f}_{4}^{2}$$\mathsf{e}_{1}$$\mathsf{e}_{2}$$\mathsf{e}_{3}$$\mathsf{e}_{4}$$\mathsf{e}_{5}$
Figure 2.1. The infinite ladder graph.
Suppose that the edge lengths of $\mathcal{G}$ are given by $\ell_{\mathsf{f}_{n}^{1}}=\ell_{\mathsf{f}_{n}^{2}}=\frac{1}{2^{n}}$ and $\ell_{\mathsf{e}_{n}}=1$ for all $n\in\mathbb{N}$. Then, the diameter of $\mathcal{G}$ is given by
$\operatorname{Diam}(\mathcal{G})=\ell_{\mathsf{e}_{1}}+\sum_{n=1}^{\infty}\ell_{\mathsf{f}_{n}^{1}}=\sum_{n=0}^{\infty}\frac{1}{2^{n}}=2.$
Now, let $(\mathcal{G}_{n})_{n\in\mathbb{N}}$ be the compact exhaustion of $\mathcal{G}$ given by
$\mathcal{G}_{n}=\bigcup_{k=1}^{n+1}\mathsf{e}_{k}\cup\bigcup_{k=1}^{n}\mathsf{f}_{k}^{1}\cup\bigcup_{k=1}^{n-1}\mathsf{f}_{k}^{2}$
(see Figure 2.2).
$\mathsf{f}_{1}^{1}$$\mathsf{e}_{1}$$\mathsf{e}_{2}$
$\mathsf{f}_{1}^{1}$$\mathsf{f}_{2}^{1}$$\mathsf{f}_{1}^{2}$$\mathsf{e}_{1}$$\mathsf{e}_{2}$$\mathsf{e}_{3}$
$\mathsf{f}_{1}^{1}$$\mathsf{f}_{2}^{1}$$\mathsf{f}_{3}^{1}$$\mathsf{f}_{1}^{2}$$\mathsf{f}_{2}^{2}$$\mathsf{e}_{1}$$\mathsf{e}_{2}$$\mathsf{e}_{3}$$\mathsf{e}_{4}$
Figure $2.$ From left to right, the graphs $\mathcal{G}_{1}$, $\mathcal{G}_{2}$ and $\mathcal{G}_{3}$.
Then, the diameter of graph $\mathcal{G}_{n}$ is given by
$\operatorname{Diam}(\mathcal{G}_{n})=\ell_{\mathsf{e}_{1}}+\sum_{k=1}^{n}\ell_{\mathsf{f}_{k}^{1}}+\ell_{\mathsf{e}_{n+1}}=2+\sum_{k=1}^{n}\frac{1}{2^{k}}.$
Consequently, we have
$\operatorname{Diam}(\mathcal{G})=2<3=\lim_{n\rightarrow\infty}\operatorname{Diam}(\mathcal{G}_{n}).$
In particular, the graph $\mathcal{G}$ and the compact exhaustion $(\mathcal{G}_{n})_{n\in\mathbb{N}}$ indeed provide an example for strict inequality in (2.4). We leave it as an open question if metric graphs exist such that the inequality in (2.4) is strict for any compact exhaustion.
@@ 2.3. Lebesgue and Sobolev spaces
We next introduce function spaces on a metric graph $\mathcal{G}$, for which embedding theorems will be discussed in Section 3.
Any metric graph $\mathcal{G}$ has both a topological and a measure theoretical structure. This immediately defines the space of continuous functions and Lebesgue spaces; we refer to [53] and [43, Sections 2.3 and 3.1] for the details of these function spaces.
@@
[…]
@@
Replacing $L^{2}$-spaces by $L^{p}$-spaces in the definition, we also obtain the Sobolev spaces $W^{1,p}(\mathcal{G})$ for $p\geq 1$; in particular, $W^{1,2}(\mathcal{G})=H^{1}(\mathcal{G})$. It is not hard to show that, for any $f\in H^{1}(\mathcal{G})$, there is a canonical extension of $f$ to the ends of $\mathcal{G}$ (see [43, Definition 3.4]). In particular, if $\overline{\mathcal{G}}$ is compact (see Corollary A.11 for a sufficient condition), then by [43, Lemma 3.2] $f\in H^{1}(\mathcal{G})$ can be identified with a function on $\overline{\mathcal{G}}$, and up to a canonical identification, we have $H^{1}(\mathcal{G})\subset C(\overline{\mathcal{G}})$.
@@
(2.7) $H^{1}_{c}(\mathcal{G}):=\left\{f\in H^{1}(\mathcal{G}):f(x)=0\text{ outside a compact subset of }\mathcal{G}\right\}$
@@
(2.8) $H^{1}_{0}(\mathcal{G}):=\overline{H^{1}_{c}(\mathcal{G})}^{\|\cdot\|_{H^{1}(\mathcal{G})}}.$
@@
    that is, $H^{1}_{0}(\mathcal{G})$-functions satisfy a Dirichlet (zero) condition at every end, see [43, Theorem 3.12].
@@
    A unified way to prescribe Dirichlet conditions at vertices, is as follows: remove $\mathsf{v}$ and appropriately add infinitely many dummy vertices on all edges incident to it. This way, $\mathsf{v}$ is replaced by $\deg(\mathsf{v})$ many ends $\gamma$, and every $f\in H^{1}_{0}(\mathcal{G})$ must vanish at $\mathsf{v}$.